Darlene built a small triangular patio on the back of her house.
The length of each of the equal and perpendicular sides of the patio is represented by the letter x. The longest side of the patio is 16 feet. What is the approximate measure , in feet, of x?
The answer is 11.3. I feel I need to use the pythagorian to solve, but I am comming up with x^2+x^2=256, and get lost there. help?

Answer by jim_thompson5910(35256)

You're on the right track

a^2+b^2 = c^2

x^2+x^2 = 16^2

x^2+x^2 = 256

2x^2 = 256

x^2 = 256/2

x^2 = 128

x = sqrt(128)

x = sqrt(64*2)

x = sqrt(64)*sqrt(2)

x = 8*sqrt(2)

x = 8*1.41421356

x = 11.313708

Round this to the nearest tenth to get x = 11.3